How much must you invest today in order to receive $10,000 at the end of each year for the next

8 years assuming you can earn 5 percent​ interest?
Answer #1

Answer : 6,768.39 (or) 6,736.40

Current investment = (10,000)*(1/1.05)^8 = 10,000*0.676839 = 6,768.39 (or) 6,768.40

r= 5%

n= 8